At the instant shown, cars A and B are traveling at speeds of 30 km/h and 20 km/h, respectively. If A is increasing its speed at 400 km/h² whereas the speed of B is decreasing at 800 km/h², determine the velocity and acceleration of B with respect to.
1
Expert's answer
2021-10-31T17:58:03-0400


Normal acceleration of B:

"(a_B)_n = \frac{v_B^2}{p} = \frac{20^2}{0.3}=1333.34 \;km/h^2 \\

Normal\; acceleration\; vector \; of \; B: \\

(a_B)_n = (a_b)_n[cos 30°i -sin 30°j] \\

= 1333.34[cos 30°i -sin 30°j] \\

= 1154.67i -666.67j"

Tangential acceleration vector of B

"(a_B)_t = (a_B)_t [-sin30°i -cos30°j] \\

= 800[-sin30°i -cos30°j] \\

= -400i -692.8j \\

Total \;acceleration \;of \;B \\

a_B = (a_B)_t +(a_B)_n \\

= -400i -692.8j +1154.67i -666.67j \\

= 754.67i -1359.47j \\

Acceleration \; of\; B\; with \; respect\; to \; A. \\

a_{B/A} = a_B -a_A \\

= (754.67i -1359.47j) -400j \\

= 754.67i -1759.47j \\

a_{B/A} = \sqrt{(754.67)^2 +(1759.47)^2} \\

= 1914.487 \;km/h^2 \\

θ = tan^{-1} (\frac{-1759.47}{754.67}) \\

θ = -66.784°"

Velocity

"v_B = -v_bcos 30°i+v_Bsin30°j \\

= -20cos30°i +20sin30°j \\

v_B = -17.32i +10j \\

v_A = -v_Ai =-30i \\

v_B =v_A +v_{B/A} \\

v_{B/A} = v_B -v_A \\

= -17.32i +10j +30i \\

= 12.68i +10j \\

= \sqrt{(12.68)^2+(10)^2} \\

= 16.148 \;km/h \\

θ = tan^{-1}(\frac{10}{12.68}) \\

θ = 38.26°"